From the ski lodge example, suppose that the probability of breaking your leg is 0.20. (Therefore the probability of not breaking your leg is (1 – 0.20) = 0.80.) What is the probability that exactly 2 out of 5 people will break their leg? n = 5, x = 2, p = .2, q = .8. What is the probability that less than 2 out of 5 people will break their leg?
Answer by stanbon(48568)

From the ski lodge example, suppose that the probability of breaking your leg is 0.20. (Therefore the probability of not breaking your leg is (1 – 0.20) = 0.80.)
What is the probability that exactly 2 out of 5 people will break their leg?
n = 5, x = 2, p = .2, q = .8.
P(x=2) = 5C2(0.2)^2*(0.8)^3 = 10*0.04*0.512 = 0.2048
------------------------------------------------------
What is the probability that less than 2 out of 5 people will break their leg?
P(x= 0 or 1) = 5C0(0.2)^0*(0.8)^5 + 5C1(0.2)^1*(0.8)^4 = 0.73728...
==============================
Cheers,
Stan H.